OpenStudy (anonymous):

Is there any way to simplify this:

OpenStudy (anonymous):

\[pir(r+\sqrt{(81/\pi(r)^2)^2+r^2}\]

OpenStudy (anonymous):

r is the radius, it's not pir!

OpenStudy (freckles):

\[\pi r(r+\sqrt{(\frac{81}{\pi r^2})^2+r^2})\]

OpenStudy (freckles):

is that right ?

OpenStudy (anonymous):

Yes! How did you get that?

OpenStudy (freckles):

you mean the code ?

OpenStudy (anonymous):

No, how did you simplify it? Oh and my notes says the answer is what you put, except there is no r+ (root)

OpenStudy (freckles):

i was just trying to copy what you had

OpenStudy (anonymous):

Ooooh my mistake, lol.

OpenStudy (freckles):

like i was just wanting to know that's what i'm/you;re suppose to simplify right?

OpenStudy (anonymous):

Yes.

OpenStudy (freckles):

well i guess you could do this: \[\pi r (r+\sqrt{(\frac{81}{\pi r^2})^2+r^2}) \\ \pi r (r+\sqrt{(\frac{1}{\pi r^2})^2} \sqrt{81^2+(\pi r^2)^2r^2})\]

OpenStudy (freckles):

then that one part the square root cancels that squared

OpenStudy (freckles):

distribute the pi*r

OpenStudy (freckles):

\[\pi r (r+\frac{1}{r \pi^2} \sqrt{81^2+\pi^2 r^6})\]

OpenStudy (anonymous):

Thank you!

OpenStudy (freckles):

\[\pi r^2 +\frac{1}{\pi} \sqrt{81^2+\pi^2 r^6}\]

OpenStudy (freckles):

but i don't know if that is really considered more simplified then what you had :p

OpenStudy (anonymous):

Eh, it's fine! It's a weird problem.